OpenStudy (anonymous):

The speed of sound V near the Earth's surface can be represented by this equation, V = 20vt+273 , where t is the temperature in degrees Celsius. Find the temperature when V = 380 meters per second. Round to the nearest whole number. A. 88 B. 90 C. 92 D. 94

OpenStudy (anonymous):

\[V = 20\sqrt{t + 273}\]

OpenStudy (anonymous):

\[V = 20\sqrt{t + 273}\] \[380=20\sqrt{t + 273}\] \[19=\sqrt{t+273}\] \[361=t+273\] \[t=88\]

Directrix (directrix):

V = 20 √ (t + 273) 380 = 20 √ (t + 273) 19 = √ (t + 273) Squaring both sides. 361 = t + 273 361 - 273 = t 88 = t

Directrix (directrix):

A. 88---> Answer